OpenStudy (anonymous):

To make 1litre of orange paint i need 700ml of red paint and 300 ml of yellow paint. How much of each will i need to make 6 litres of orange paint? (plz help!!)

OpenStudy (anonymous):

700*6=4200ml ...300*6=1800ml

OpenStudy (anonymous):

orange paint = 6 * 700 ml = 4200 ml or 4 l and 200 ml yellow paint = 6 * 300 ml = 1800 ml or 1 l and 800 ml -------- 6000 ml i.e. 6 litres

OpenStudy (anonymous):

thankkk you ur a great help!